OpenStudy (anonymous):

Kapil Dev and Bishan Singh Bedi started running at 8 am with a speed of 40 km/hr and 18 km/hr respectively for Pragati Maidan from the India gate. Bedi completes the whole distance with a uniform speed. After 1 ½ hr, Kapil Dev was stuck in traffic and reduced his speed. In the next 30 minutes, he could cover only 15 km and covered the remaining distance with a speed equal in magnitude to the total time taken by Bedi in hours to cover the entire journey and reached Pragati Maidan at the same time. At what time did both of them reach Pragati Maidan?

OpenStudy (anonymous):

@anteater

OpenStudy (anonymous):

hi ,after three weeks i saw u in os

OpenStudy (anteater):

Hello :)

OpenStudy (anteater):

Busy summer!

OpenStudy (anonymous):

60km -kapil without struck

OpenStudy (anonymous):

for 2 hours he can cover only 75 km

OpenStudy (anteater):

Yes, so when he isn't stuck he covers 60 km

OpenStudy (anteater):

I think I need to work it out on paper first :)

OpenStudy (anonymous):

and remaining equal to =18*2=36 36+x

OpenStudy (anteater):

That sounds right, but let me double check

OpenStudy (anonymous):

75+(40-x)y=36+36*y

OpenStudy (anonymous):

y is time ,x is reduced speed

OpenStudy (anonymous):

so sorry 75+(40-x)y=36+18*y

OpenStudy (anteater):

That's ok :) I was trying to leave it all in terms of x in my calculations. I am looking at it again!

OpenStudy (anonymous):

oh sry still i made a mistake 75+(40-x)y=18*y

OpenStudy (anteater):

That is ok! I think you are on the right track.

OpenStudy (anonymous):

75-xy+22y=0

OpenStudy (anteater):

Good! Now we just need to get x in terms of y (or vice versa)

OpenStudy (anonymous):

y*(22-x)=-75

OpenStudy (anonymous):

i think x=22

OpenStudy (anteater):

So y represents the time it took Bedi to run the race. And x is Dev's reduced speed?

OpenStudy (anonymous):

yes

OpenStudy (anteater):

Ok, thanks! Just a moment, please.

OpenStudy (anonymous):

k.yw

OpenStudy (anteater):

I think you are on the right track, but let's look at the last part where Dev slows his pace again.

OpenStudy (anteater):

You have the total distance is 18y, where y is the amount of time it takes Bedi to run the race. And since they both started at 8 am and reached the end of the race at the same time, they both have to have been running for y hours total.

OpenStudy (anonymous):

yes already 2 hours over for dev

OpenStudy (anteater):

Yes :) Mind if I draw? It makes it easier for me to keep track.

OpenStudy (anteater):



OpenStudy (anteater):

I think it would work out that way, since Bedi runs 18y miles, where y is the number of hours he runs. Then Dev's distance is broken into parts. He runs for 1.5 hours at 40 km/h, then in the next .5 hour he covers 15 km, and then in the remaining time (y - 2 hours) he runs at a reduced speed, x.

OpenStudy (anteater):

And then they tell you that the reduced speed is equal to the total number of hours that Bedi ran. So x = y

OpenStudy (anonymous):

they said the distance covered with a speed equal to time taken by bedi so it will be y not y-2.

OpenStudy (anteater):

Yes, but he is running at a rate of y for the length of time y-2

OpenStudy (anteater):

Would you agree that they both ran for y hours total?

OpenStudy (anteater):

Since they both left at the same time and arrived at the same time?

OpenStudy (anonymous):

yes

OpenStudy (anteater):

So then, for Dev, the first part of his race he runs for 1.5 hours (at 40 km/h); the second part of the race he runs for .5 hour (and they tell you he covers 15 km during that part). And so, how long must he run in the final part of the race, the part where he is running at a reduced speed?

OpenStudy (anonymous):

20y-xy-5=0

OpenStudy (anteater):

So if he runs for a total of y hours, and he has already (as you said) run for 2 hours, the remaining time would be ...

OpenStudy (anonymous):

y(20-x)=5

OpenStudy (anteater):

Let me check :)

OpenStudy (anonymous):

x=20,y=5

OpenStudy (anteater):

So if Bedi ran for 5 hours, it would be a distance of 90 km (since 18 X 5 = 90)

OpenStudy (anonymous):

yes dev will be135

OpenStudy (anteater):

Then, Dev has already covered 75 km. 90 - 75 = 15 km

OpenStudy (anteater):

15 km is the remaining distance, and if Dev covers this at a rate of 5 km/hour, it will take him 3 hours to cover that part of the race. He has already run 2 hours (1.5 + .5 = 2), so then tacking on another 3 hours will mean that he has run 5 hours as well. Good!!!! :D

OpenStudy (anteater):

So ... we now know that they have both run for 5 hours. If they started at 8 am, what time will it be?

OpenStudy (anonymous):

1 pm

OpenStudy (anteater):

Yup!

OpenStudy (anteater):

After which they both collapsed of heat stroke.

OpenStudy (anteater):

It occurs to me that there could be an additional "correct" answer to this problem. If they both ran for 15 hours, they would cover a distance of 270 km. In the first part of the race, Dev would cover 60 km, in the second he would cover 15 km, and then in the final part he would cover 15(15-2) or 195 km. So, if they both ran 15 hours they would arrive at 11 pm. But that seems a bit unrealistic for a run! Ultramarathoners :) A 5-hour run seems more likely. 

OpenStudy (anonymous):

yupp:)

OpenStudy (anteater):

Just a moment ... going to find out how long the longest ultramarathon has been ...

OpenStudy (anteater):

I guess that would be this one: http://en.wikipedia.org/wiki/Self-Transcendence_3100_Mile_Race

OpenStudy (anteater):

I think your feet might hate you after that!

OpenStudy (anonymous):

oh ur telling the race! right:-)

OpenStudy (anonymous):

yessssssssssss........................

OpenStudy (anteater):

LOL !

OpenStudy (anonymous):

can we do another problem

OpenStudy (anteater):

Sure!

OpenStudy (anonymous):

lol

OpenStudy (anteater):

Did you post another problem on the board? I apologize if I have missed it somehow!

OpenStudy (anonymous):

yes i tagged it

OpenStudy (anteater):

Oh! Ok ... Just a moment, then!